OpenStudy (anonymous):

under what conditions of temperature and pressure would you expect the behaviour of carbon dioxide to be less ideal than of hydrogen

OpenStudy (aaronq):

what are the conditions at which gases behave most ideally?

OpenStudy (anonymous):

at low temperature and high pressure carbon dioxide behave less ideal.

OpenStudy (aaronq):

its the other way around, high temp and low pressure

OpenStudy (theeric):

"The ideal gas model tends to fail at lower temperatures or higher pressures..." - http://en.wikipedia.org/wiki/Ideal_gas

OpenStudy (theeric):

The full sentence is: "The ideal gas model tends to fail at lower temperatures or higher pressures, when intermolecular forces and molecular size become important."

OpenStudy (theeric):

I might be too tired to think well, but I think I put a few things together. -The ideal gas [state] is one in which the gas molecules hardly interact due to molecular characteristics. Thus, the pressure, volume, and temperature relationship expressed as \(P\ V=n\ R\ T\) in the Ideal Gas Law is fairly accurate. That is, the pressure, volume, and temperature are the more significant factors of gas molecule interaction. Now, I guess that a low temperature would mean that the molecules would have lower energies and thus molecular forces like electromagnetic forces matter more than the slight randomness given by slight thermal energy. And I also guess that high pressure would mean the gas molecules are tightly pushed together, and are more likely to be close in proximity. In close proximity, electromagnetic forces are strong. But they decay with the increased square of the distance between them. So, you have two objects 3m apart, and move them so they're 5m apart. The change is distance is 2m, but the change in the squares of the distances are \(5^2-3^2=25-9=16\). Move them two more meters apart, from 5m to 7m, and the change in the squares of the distances is \(7^2-5^2=49-25=24\). As you go farther away, the strength of the electric forces decay quicker. For gas molecules to get in closer proximity to each other makes a huge difference on how they interact. I hope this helps, and I hope somebody can verify my thoughts.

OpenStudy (aaronq):

"at low temperature and high pressure carbon dioxide behave less ideal" sorry, i read that as "behave most ideal". you're right, mah bad.

OpenStudy (theeric):

Haha, Wikipedia is great.